Shift Summary for Owl, 8/25/99

Ed Blucher (blucher@hep.uchicago.edu)
Wed, 25 Aug 1999 07:05:22 -0500


Shift Summary for Owl Shift, 08/25/99

On Shift: Ed Blucher, Bob Hsiung

Hours of beam available at all intensities: 4.9
Hours of beam available at nominal intensity: 4.9
Hours of E832 data (special runs not included): 4.7
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: TeV quench, MI probs, kfrends

========================================================================
Runs ended this shift:
Run 14322 - BEAM_832 (in progress)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking (SEM=6.8E12)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14322 2:40 TeV quench at A1; estimate 45 minutes

4:35 Beam returns

278 5:09 DYC full, crate 53 (AZLAT)

292 5:28 DYC full, laser crate

307 5:50 DYC full, crate 53 (AZLAT)

310 5:53 No beam. Main injector beam fell out. Investigating.

321 6:07 Beam returns.

6:50 Mount tape QKA018 for onsplit.

6:58 warning from onsplit: tape ktev3_tps2d4 (tape QKA018)
is in UNUSED_ALLOCATED state. (The tape is being
written)
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Check with Sydney about onsplit warning described above.